Solve by substitution.
x - y = 3
3x - 2y = 3

Answer by dead_famous1991(7)

Solve by substitution.

x - y = 3
3x - 2y = 3

x - y = 3
=>x = 3 + y

3x - 2y = 3
=>3x = 3 + 2y
=>3 (3+y) = 3 + 2y
=>9 + 3y = 3 + 2y
=>3y - 2y = 3 - 9
=>y = -6

x - y = 3
=>x - (-6) = 3
=>x + 6 = 3
=>x = 3 - 6
=>x = -3